# PID Accelerated Temporal Difference Algorithms

### v1

@@ B.1 Proof of Theorem 2
Suppose synchronous TD Learning and PID TD Learning are run with initial value function $V_{0}$ and learning rate $\mu(t)=\epsilon/(t+T)$ to evaluate policy $\pi$. Let $V^{\mathrm{TD}}_{t},V^{\mathrm{PID}}_{t}$ be the value function obtained with each algorithm at iteration $t$. Assume $\delta^{\mathrm{TD}},\delta^{\mathrm{PID}}\geq 0$. If ${\mathcal{P}}^{\pi}$ is not diagonalizable, we further assume $\delta^{\mathrm{TD}}>0$, and if $A_{g}^{\pi}$ is not diagonalizable, we assume $\delta^{\mathrm{PID}}>0$. If $\epsilon>2/(1-\gamma-\delta^{\mathrm{TD}})$ and $T\geq c^{\mathrm{PID}}_{1}\epsilon/(1-\gamma-\delta^{\mathrm{TD}})$, we have
@@
Here, $\{c^{\mathrm{PID}}_{i}\}$ are constants dependent on the MDP and $\delta^{\mathrm{TD}}$. Moreover, assume we initialize $V^{\prime}=V_{0},z\equiv 0$ in PID TD Learning and $A_{g}^{\pi}$ has spectral radius $\rho<1$. If $\epsilon>2/(1-\rho-\delta^{\mathrm{PID}})$ and $T\geq c^{\mathrm{PID}}_{1}\epsilon/(1-\rho-\delta)$, we have
@@
which means $T^{\pi}$ is a $(\gamma+\delta^{\mathrm{TD}})$-contraction w.r.t $\left\lVert\cdot\right\rVert_{2,S}$. Moreover, $\frac{1}{2}\left\lVert x\right\rVert_{2,S}^{2}$ is $1$-smooth according to Lemma 3. Consequently, we use $\left\lVert\cdot\right\rVert_{2,S}$ as the norms $\left\lVert\cdot\right\rVert_{c}$ and $\left\lVert\cdot\right\rVert_{s}$ in (Chen et al., 2020b).
Assume the policy is $d$-deterministic. By Lemma 4, we can bound the noise at each iteration as
$\mathbb{E}\left[\left\lVert W_{t}\right\rVert_{\infty}^{2}\right]\leq C^{\mathrm{TD}}+B^{\mathrm{TD}}\left\lVert\tilde{V}_{t}\right\rVert^{2}_{\infty},$
@@
By Corollary 2.1.2 of Chen et al. (2020b), choosing the norm $\left\lVert\cdot\right\rVert_{e}$ to be $\left\lVert\cdot\right\rVert_{\infty}$, the error at iteration $t$ is bounded as
@@
which means $L_{g}^{\pi}$ is a $(\rho+\delta^{\mathrm{PID}})$-contraction w.r.t $\left\lVert\cdot\right\rVert_{2,S}$. Moreover, $\frac{1}{2}\left\lVert x\right\rVert_{2,S}^{2}$ is $1$-smooth according to Lemma 3. Consequently, we use $\left\lVert\cdot\right\rVert_{2,S}$ as the norms $\left\lVert\cdot\right\rVert_{c}$ and $\left\lVert\cdot\right\rVert_{s}$ in (Chen et al., 2020b).
Assume the policy is $d$-deterministic. By Lemma 4, we can bound the noise at each iteration as
$\mathbb{E}\left[\left\lVert\tilde{W}_{t}\right\rVert_{\infty}^{2}\right]\leq C^{\mathrm{PID}}+B^{\mathrm{PID}}\left\lVert\tilde{V}_{t}\right\rVert^{2}_{\infty},$
@@
By Corollary 2.1.2 of Chen et al. (2020b), choosing the norm $\left\lVert\cdot\right\rVert_{e}$ to be $\left\lVert\cdot\right\rVert_{\infty}$, the error at iteration $t$ is bounded as
@@ D.1 Chain Walk
The environment consists of 50 states that are connected in a circular chain. The agent has two actions available, moving left or right. Upon taking an action, the agent succeeds with probability 0.7, stays in place with probability 0.1, and moves in the opposite direction with probability 0.2. The agent receives a reward of 1 when entering state 10, and reward -1 when entering state 40, and a reward of 0 otherwise.
The policy evaluated in the TD experiments is to always move left.
@@ D.2 Cliff Walk
A 6 by 6 grid world is used, visualized in Figure 5. The agent starts on the top left its goal is to end up on the top right. There are 12 cliff tiles, and the agent is stuck in them if it falls in. Moreover, the agent is stuck in the goal state once entering it. Upon making a move in the goal state, it receives a reward of 20. Making a move in a cliff receives a reward of -32, -16, or -8 depending on if the cliff is at the top, middle, or bottom respectively. Otherwise, it receives a reward of -1.
The agent has four possible actions corresponding to moving up, down, left, and right. If the agent attempts to move off the grid, it simply stays in place. Otherwise, its action succeeds with probability 0.9, and moves in one of the other three directions at random with uniform probability
The policy evaluated in the TD experiments is a random walk.
@@ D.3 Garnet
The environment is randomly generated. They consist of 50 states, and 3 actions per state.
To build the environment, for each action and state, $(x,a)$, we pick 5 other random states ${\mathcal{X}}_{x,a}$. For 10 randomly chosen states $x$, we set $r(x)$ from a uniform distribution between 0 and 1. We set $r(x)$ is zero on all other states. Then, when taking action $a$ and from state $x$, we receive reward $r(x)$ and move to any state in ${\mathcal{X}}_{x,a}$ with equal probability.
The policy evaluated in the TD experiments is a random walk.
@@ Appendix E Additional Experiments
Figure 6: PID TD Learning with Gain Adaptation in Chain Walk with $\gamma=0.99$. (Left) Comparison of value errors of PID TD Learning with TD Learning. Each curve is average over 80 runs. Shaded area shows standard error. (Right) The change of gains done by Gain Adaptation through training.
Figure 7: Q-Learning with Gain Adaptation in Cliff Walk with $\gamma=0.99$. (Left) Comparison of value errors of PID Q-Learning with Q-Learning. Each curve is average over 80 runs. Shaded area shows standard error. (Right) The change of gains done by Gain Adaptation through training.
@@ Appendix F Details of Experimental Setup
We pick the hyperparameters such that a normalized error of $0.2$ is achieved the fastest, and if this error is not achieved, the final error is minimized.
We fix $\alpha=0.05$, $\beta=0.95$, and $\lambda=0.5$ throughout all the experiments. For the Cliff Walk policy evaluation experiments in Figure 2, we set $\eta=10^{-5}$ and $\epsilon=0.1$. For the Chain Walk (Control) experiments in Figure 3, we set $\eta=4\times 10^{-8}$ and $\epsilon=10^{-4}$. For the Chain Walk (PE) experiments in Figure 6, we set $\eta=5\times 10^{-7}$ and $\epsilon=10^{-3}$. For the Cliff Walk (Control) experiments in Figure 7, we set $\eta=8\times 10^{-9}$ and $\epsilon=10^{-3}$.
For the Garnet (PE) experiments in Figure 4, we perform a grid search on $\eta\in\{0.1,0.01,0.001,0.0001\}$, $\epsilon\in\{0.1,0.01\}$. Similarly, for the Garnet (Control) experiments, we use $\epsilon=0.1$ and perform a grid search is over $\eta\in\{10^{-5},5\times 10^{-5},10^{-6}\}$. The learning rates we perform a grid search over in these tests is listed in Table 3. The grid search is separately performed for each instance of the sampled Garnet.
On each randomly generated Garnet environment, 80 runs are performed and the average trajectory is found. The variation of this average trajectory among all the 80 Garnet environments is shaded in Figure 4.
Learning Rates ($\epsilon:\text{associated set of }M$)
@@ D Component
Table 3: List of learning rates searched through in the Garnet experiments.

### v2

@@ B.1 Proof of Theorem 2
Suppose synchronous TD Learning and PID TD Learning are run with initial value function $V_{0}$ and learning rate $\mu(t)=\epsilon/(t+T)$ to evaluate policy $\pi$. Let $V^{\mathrm{TD}}_{t},V^{\mathrm{PID}}_{t}$ be the value functions obtained with each algorithm at iteration $t$. Assume $\delta^{\mathrm{TD}},\delta^{\mathrm{PID}}\geq 0$. If ${\mathcal{P}}^{\pi}$ is not diagonalizable, we further assume $\delta^{\mathrm{TD}}>0$, and if $A_{g}^{\pi}$ is not diagonalizable, we assume $\delta^{\mathrm{PID}}>0$. If $\epsilon>2/(1-\gamma-\delta^{\mathrm{TD}})$ and $T\geq c^{\mathrm{PID}}_{1}\epsilon/(1-\gamma-\delta^{\mathrm{TD}})$, we have
@@
Here, $\{c^{\mathrm{PID}}_{i}\}$ are constants dependent on the MDP and $\delta^{\mathrm{TD}}$. Moreover, assume we initialize $V^{\prime}=V_{0},z\equiv 0$ in PID TD Learning and $A_{g}^{\pi}$ has spectral radius $\rho<1$. If $\epsilon>2/(1-\rho-\delta^{\mathrm{PID}})$ and $T\geq c^{\mathrm{PID}}_{1}\epsilon/(1-\rho-\delta^{\mathrm{PID}})$, we have
@@
which means $T^{\pi}$ is a $(\gamma+\delta^{\mathrm{TD}})$-contraction w.r.t. $\left\lVert\cdot\right\rVert_{2,S}$. Moreover, $\frac{1}{2}\left\lVert x\right\rVert_{2,S}^{2}$ is $1$-smooth according to Lemma 3. Consequently, we use $\left\lVert\cdot\right\rVert_{2,S}$ as the norms $\left\lVert\cdot\right\rVert_{c}$ and $\left\lVert\cdot\right\rVert_{s}$ in Chen et al. (2020).
Assume the policy is $d$-deterministic. Define the noise $W_{t}\colon{\mathcal{X}}\to\mathbb{R}$ at iteration $t$ as
$W_{t}(x)\triangleq R_{x,t}+\gamma V_{t}(X^{\prime}_{x,t})-r^{\pi}(x)-\gamma({\mathcal{P}}^{\pi}V_{t})(x).$
By Lemma 4, we can bound the conditional noise at each iteration as
$\mathbb{E}\left[\left\lVert W_{t}\right\rVert_{\infty}^{2}\middle|V_{0},W_{0},\ldots,W_{t-1},V_{t}\right]\leq C^{\mathrm{TD}}+B^{\mathrm{TD}}\left\lVert V_{t}\right\rVert^{2}_{\infty},$
@@
By Corollary 2.1.2 of Chen et al. (2020), choosing the norm $\left\lVert\cdot\right\rVert_{e}$ to be $\left\lVert\cdot\right\rVert_{\infty}$, $\mu=L=1$, the error at iteration $t$ is bounded as
@@
which means $L_{g}^{\pi}$ is a $(\rho+\delta^{\mathrm{PID}})$-contraction w.r.t. $\left\lVert\cdot\right\rVert_{2,S}$. Moreover, $\frac{1}{2}\left\lVert x\right\rVert_{2,S}^{2}$ is $1$-smooth according to Lemma 3. Consequently, we use $\left\lVert\cdot\right\rVert_{2,S}$ as the norms $\left\lVert\cdot\right\rVert_{c}$ and $\left\lVert\cdot\right\rVert_{s}$ in Chen et al. (2020).
Assume the policy is $d$-deterministic. Define the noise $\tilde{W}_{t}\colon{\mathcal{X}}\times\{\texttt{v},\texttt{z},\texttt{v'}\}\to\mathbb{R}$ at iteration $t$ as
$\tilde{W}_{t}(x,\texttt{v})$ $\triangleq(\kappa_{p}+\kappa_{I}\alpha)(R_{x,t}+\gamma V_{t}(X^{\prime}_{x,t})-r^{\pi}(x)-\gamma({\mathcal{P}}^{\pi}V_{t})(x)),$
$\tilde{W}_{t}(x,\texttt{z})$ $\triangleq\alpha(R_{x,t}+\gamma V_{t}(X^{\prime}_{x,t})-r^{\pi}(x)-\gamma({\mathcal{P}}^{\pi}V_{t})(x)),$
$\tilde{W}_{t}(x,\texttt{v'})$’ ) $\triangleq 0.$
By Lemma 4, we can bound the noise at each iteration as
$\mathbb{E}\left[\left\lVert\tilde{W}_{t}\right\rVert_{\infty}^{2}\middle|\tilde{V}_{0},\tilde{W}_{0},\ldots,\tilde{W}_{t-1},\tilde{V}_{t}\right]\leq C^{\mathrm{PID}}+B^{\mathrm{PID}}\left\lVert\tilde{V}_{t}\right\rVert^{2}_{\infty},$
@@
By Corollary 2.1.2 of Chen et al. (2020), choosing the norm $\left\lVert\cdot\right\rVert_{e}$ to be $\left\lVert\cdot\right\rVert_{\infty}$, the error at iteration $t$ is bounded as
@@ D.1 Chain Walk
The environment consists of 50 states that are connected in a circular chain. The agent has two actions available, moving left or right. Upon taking an action, the agent succeeds with probability 0.7, stays in place with probability 0.1, and moves in the opposite direction with probability 0.2. The agent receives a reward of 1 when entering state 10, a reward of -1 when entering state 40, and a reward of 0 otherwise. The policy evaluated in the PE experiments is to always move left.
@@ D.2 Cliff Walk
A 6 by 6 grid world is used, visualized in Figure 5. The agent starts on the top left. Its goal is to end up on the top right. There are 12 cliff tiles, and the agent is stuck in them if it falls in. Moreover, the agent is stuck in the goal state once entering it. Upon making a move in the goal state, it receives a reward of 20. Making a move in a cliff receives a reward of -32, -16, or -8 depending on whether the cliff is on the top, middle, or bottom respectively. Otherwise, it receives a reward of -1. The agent has four possible actions corresponding to moving up, down, left, and right. If the agent attempts to move off the grid, it simply stays in place. Otherwise, its action succeeds with probability 0.9, and moves in one of the other three directions at random with uniform probability The policy evaluated in the TD experiments is a random walk.
@@ D.3 Garnet
The environment is randomly generated. They consist of 50 states, and 3 actions per state. To build the environment, for each action and state, $(x,a)$, we pick 5 other random states ${\mathcal{X}}_{x,a}$. For 10 randomly chosen states $x$, we set $r(x)$ from a uniform distribution between 0 and 1. We set $r(x)$ is zero on all other states. Then, when taking action $a$ and from state $x$, we receive reward $r(x)$ and move to any state in ${\mathcal{X}}_{x,a}$ with equal probability. The policy evaluated in the TD experiments is a random walk.
@@ Appendix E Additional Experiments
Figure 6: PID TD Learning with Gain Adaptation in Chain Walk with $\gamma=0.99$. (Left) Comparison of value errors of PID TD Learning with TD Learning. Each curve is averaged over 80 runs. Shaded area shows standard error. (Right) The change of gains done by Gain Adaptation through training.
Figure 7: Q-Learning with Gain Adaptation in Cliff Walk with $\gamma=0.99$. (Left) Comparison of value errors of PID Q-Learning with Q-Learning. Each curve is averaged over 80 runs. Shaded area shows standard error. (Right) The change of gains done by Gain Adaptation through training.
@@ Appendix F Details of Experimental Setup
We pick the hyperparameters such that a normalized error of $0.2$ is achieved the fastest, and if this error is not achieved, the final error is minimized. We fix $\alpha=0.05$, $\beta=0.95$, and $\lambda=0.5$ throughout all the experiments.
For the Garnet (PE) experiments in Figure 4, we perform a grid search on $\eta\in\{0.1,0.01,0.001,0.0001\}$, $\epsilon\in\{0.1,0.01\}$. Similarly, for the Garnet (Control) experiments, we use $\epsilon=0.1$ and perform a grid search over $\eta\in\{10^{-5},5\times 10^{-5},10^{-6}\}$. The learning rates we perform a grid search over in these tests is listed in Table 3. The grid search is separately performed for each instance of the sampled Garnet. For TD Learning and Q-learning, the rates searched over are the same as that of the P component in Table 3. On each randomly generated Garnet environment, 80 runs are performed and the average trajectory is found. The variation of this average trajectory among all the 80 Garnet environments is shaded in Figure 4.
For the Cliff Walk policy evaluation experiments in Figure 2, we set $\eta=10^{-5}$ and $\epsilon=10^{-1}$. For the Chain Walk (Control) experiments in Figure 3, we set $\eta=4\times 10^{-8}$ and $\epsilon=10^{-4}$. For the Chain Walk (PE) experiments in Figure 6, we set $\eta=5\times 10^{-7}$ and $\epsilon=10^{-1}$. For the Cliff Walk (Control) experiments in Figure 7, we set $\eta=10^{-8}$ and $\epsilon=10^{-1}$. For picking the learning rate for the I and D component, we also consider learning rates of the form $\min(0.25,N_{t}(X_{t})/M)$ with $M\in\{\infty,10,100,500,1000,10000\}$.
Learning Rates $\min(\epsilon,N_{t}(X_{t})/M)$ searched through
(formatted $\epsilon:\text{corresponding set of }M$)
@@ D Component
Table 3: All the learning rates searched through in the Garnet experiments.